OpenStudy (anonymous):

Can someone simplify this completely? show steps please? (√2- 4i)(√2 +4i)

myininaya (myininaya):

(A+B)(A-B)=A^2-B^2 so we have \[(\sqrt{2}-4i)(\sqrt{2}+4i)=(\sqrt{2})^2-(4i)^2\]

myininaya (myininaya):

\[=2-4^2i^2=2-(16)(-1)=2+16=18\]

OpenStudy (anonymous):

just foil. multiple first set of numbers sqrt2* sqrt2=2 then outer numbers sqrt2*4i =4isqrt2 then inner number qrt2*4i =4isqrt2 then last 4i*4i=16 i then answer is =2+4i*sqrt2-41*sqrt2-16i which when simplified = 2+16i i think.....

OpenStudy (anonymous):

wait actually it'll be 2+16(i^2) = 2+16 = 18

OpenStudy (anonymous):

aaaaaaaaaaaaaaaaarghhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

\[(a+bi)(a-bi)=a^2+b^2\] finished

OpenStudy (anonymous):

\[\sqrt{2}^2=2\] \[4^2=16\] \[2+16=18\]

OpenStudy (anonymous):

thank you everyone. my book has no example and my teacher isn't that good at explaining.

OpenStudy (anonymous):

forget the "foil" and forget the flippin' "i"just know that a complex number multiplied by its complex conjugate is the sum of the squares

OpenStudy (anonymous):

but it's important that she does it the way she is learning in school, even if it's the longer... more complicated way... which can suck